OpenStudy (anonymous):

The formula for the amount with simple interest is A = P(1 + rt). Find t if A = 168, P = 150, and r = 0.08.

OpenStudy (mathmate):

The question is assumed to be :find t. A = P(1 + rt) Isolate t and solve: (1+rt)=P/A rt=(P/A)-1 t=[(P/A)-1]/r Substitute the given values and evaluate.

OpenStudy (anonymous):

ok ty

OpenStudy (mathmate):

You're welcome!